# Network Slicing Resource Management in Uplink User-Centric Cell-Free Massive MIMO Systems

### v1

@@ Abstract
This paper addresses the joint optimization of per-user equipment (UE) bandwidth allocation and UE–access point (AP) association to maximize weighted sum-rate while satisfying heterogeneous quality-of-service (QoS) requirements across enhanced mobile broadband (eMBB) and ultra-reliable low-latency communication (URLLC) slices in the uplink of a network slicing-enabled user-centric cell-free (CF) massive multiple-input multiple-output (mMIMO) system. The formulated problem is NP-hard, rendering global optimality computationally intractable. To address this challenge, it is decomposed into two sub-problems, each solved by a computationally efficient heuristic scheme, and jointly optimized through an alternating optimization framework. We then propose (i) a bandwidth allocation scheme that balances UE priority, spectral efficiency, and minimum bandwidth demand under limited resources to ensure fair QoS distribution, and (ii) a priority‑based UE-AP association assignment approach that balances UE service quality with system capacity constraints. Together, these approaches provide a practical and computationally efficient solution for resource‑constrained network slicing scenarios, where QoS feasibility is often violated under dense deployments and limited bandwidth, necessitating graceful degradation and fair QoS preservation rather than solely maximizing the aggregate sum‑rate. Simulation results demonstrate that the proposed scheme achieves up to $52\%$ higher weighted sum-rate, $140\%$ and $58\%$ higher QoS success rates for eMBB and URLLC slices, respectively, while reducing runtime by up to $97\%$ compared to considered benchmarks.
@@ I Introduction
The sixth-generation (6G) wireless systems must support heterogeneous services with stringent quality-of-service (QoS) demands, primarily enhanced mobile broadband (eMBB) (high-rate, reliable) and ultra-reliable low-latency communication (URLLC) (ultra-low latency, high reliability) [8]. Network slicing (NS) addresses this by partitioning physical infrastructure into logical slices, each optimized for specific QoS objectives [10]. Cell-free (CF) massive multiple-input multiple-output (mMIMO), a promising 6G architecture [11], eliminates cell boundaries by enabling distributed access points (APs) to cooperatively serve user equipments (UEs), achieving superior spectral efficiency (SE), uniform experience, and enhanced reliability through spatial diversity and user-centric (UC) connectivity [5]. However, integrating NS into CF mMIMO introduces challenges in resource allocation, UE association, and inter-slice isolation [16, 4].
@@
In the CF mMIMO systems, existing research has predominantly focused on downlink (DL) slicing. For instance, [15] proposes multi-tenant resource allocation maximizing infrastructure revenue while preserving service quality, though without dynamic link selection critical for scalability. The work in [16] introduces a hierarchical slicing framework leveraging multiple time scales to maximize cross-service utility, while [4] exploits spatial diversity to meet heterogeneous URLLC and eMBB requirements under imperfect CSI. However, they avoid the resource allocation in infeasible resource constraint cases. While uplink (UL) slicing is equally important and poses distinct challenges due to diverse UE demands and limited inter-UE coordination, it has received comparatively less attention in the literature. Nevertheless, UL slicing has not yet been explored in the context of the CF mMIMO systems.
@@
Designing UL access is inherently more complex than DL due to lack of UE coordination and different resource priorities. While DL prioritizes bandwidth and throughput for content delivery, UL emphasizes power management and timely transmission for latency-critical applications. Thus, the stringent QoS requirements imposed by emerging UL-intensive applications (autonomous vehicles, telemedicine, AR/VR) cannot be met by reusing DL slices, necessitating dedicated UL slicing mechanisms for CF mMIMO systems. However, practical high-density deployments with limited bandwidth frequently violate resource allocation feasibility, where total minimum bandwidth demand exceeds available capacity (further discussed in Section IV-A). In such scenarios, maintaining service continuity for all UEs while ensuring fair QoS distribution becomes critical to prevent complete service outage. This motivates the development of resource allocation strategies that gracefully degrade service quality under scarcity while preserving system-wide fairness and reliability.
@@
In this work, we consider UL slicing in a NS-enabled CF mMIMO framework to support heterogeneous eMBB and URLLC traffic while improving system performance while considering practical assumption like pilot contamination (PC). To the best of our knowledge, the UL slicing problem in the context of the CF mMIMO systems has not yet been addressed in the literature. The main contributions are: (i) we formulate a joint optimization for per-UE bandwidth allocation and UE–AP association to maximize weighted sum-rate in NS-enabled UL CF mMIMO systems; due to the NP-hardness of the problem, we propose an alternating-optimization (AO) approach that decomposes it into two tractable sub-problems solved iteratively, (ii) we develop two computationally efficient heuristic algorithms: an efficiency‑oriented bandwidth allocation scheme and a priority‑based UE–AP association scheme, ensuring fair QoS distribution even when feasibility is violated, thereby enabling graceful degradation under resource scarcity, and (iii) extensive numerical results demonstrate that the proposed solution achieves a favorable trade-off between weighted sum-rate and QoS satisfaction while substantially lowering computational cost compared to an interior-point solver-based solution.
@@ II-C Channel Estimation Model
The UL transmission consists of two main stages: training and data transmission. In each coherence interval of length $\tau_{c}$ symbols, $\tau_{p}$ symbols are allocated for channel estimation and the rest is for data transmission. During UL training, all UEs transmit their assigned pilot sequences, which are received by the APs for channel estimation. As in practical systems, we assume that $K\gg\tau_{p}$, implying that some UEs reuse the same pilot sequences, leading to PC. The channel is estimated using minimum mean square error (MMSE) approach and the quality of the estimated channel $\hat{\mathbf{g}}_{k,m}$ is quantified by the mean-square of the estimate: $\gamma_{k,m}\triangleq\mathbb{E}\left[\|\hat{\mathbf{g}}_{k,m}\|^{2}\right]=\sqrt{\tau_{p}\rho_{p}\,\eta_{k}^{p}\,}\beta_{k,m}\,c_{k,m}$. The MMSE scaling coefficient $c_{k,m}$ is given by
@@ II-D Uplink Data Transmission Model
During UL transmission, each UE $k\in\mathcal{K}$ sends data symbols to its cooperating APs, [^1] Each AP receives all UEs’ signals but decodes only those from its associated UEs according to the UE-AP association scheme. which forward them to the CPU for joint decoding using channel estimates. URLLC UEs, having short packets, are modeled using finite block length theory rather than the Shannon capacity [16]. The achievable rate expression for both eMBB and URLLC UEs is given at the top of the next page, where $V_{k}=1-(1+\text{SINR}_{k})^{-2}$ is the channel dispersion, ${Q}^{-1}(\cdot)$ is the inverse Gaussian Q-function, $\theta$ is the decoding error probability, and $L_{k}$ is the packet size of UE $k$. The closed-form signal-to-interference-plus-noise ratio (SINR) expression for UE $k$ ($\text{SINR}_{k}$) is written as (5) based on the similar steps described in [6], where $\eta_{k}^{d}\in(0,1]$ denotes the data power control coefficient of UE $k$ and $\rho_{d}$ is the normalized UL data SNR.
@@ II-E Delay Model
To capture the latency characteristics of delay-sensitive services such as URLLC, we adopt an $M/M/1$ queuing model as used in [14] to characterize the UL queuing and transmission delay experienced by each UE. For UE $k\in\mathcal{K}_{\mathrm{URLLC}}$, let $\lambda_{k}$ denote the average packet arrival rate (in packets per second). The service rate of UE $k$ is $\mu_{k}=R_{k}/L_{k}$ (in packets per second). So, under the stability condition $\mu_{k}>\lambda_{k}$ [7], the total delay experienced by UE $k$, denoted by $D_{k}$, comprises both queuing and transmission delays and is expressed as: $D_{k}=\frac{1}{\mu_{k}-\lambda_{k}}=\frac{1}{\frac{R_{k}}{L_{k}}-\lambda_{k}},\quad\forall k\in\mathcal{K}_{\mathrm{URLLC}}$.
To satisfy the QoS requirements of latency-critical applications, the total delay must not exceed a predefined marhomum tolerable latency threshold $D_{k}^{\mathrm{max}}$. This requirement imposes the following constraint on the achievable UL rate: $\frac{R_{k}}{L_{k}}\geq\lambda_{k}+\frac{1}{D_{k}^{\mathrm{max}}},\quad\forall k\in\mathcal{K}_{\mathrm{URLLC}}$. This constraint ensures that the combined queuing and transmission delays remain within acceptable bounds for each UE of $\mathcal{K}_{\mathrm{URLLC}}$, thereby preserving the reliability and responsiveness required by mission-critical services.
@@ III Problem Formulation
Our objective is to jointly optimize the per-UE bandwidth allocation and UE-AP association assignment for maximizing the weighted sum-rate performance under fixed slice-wise bandwidth $B_{s},\ s\in\mathcal{S}$, while satisfying service-specific QoS requirements. The weights, $w_{k}$ reflect heterogeneous service priorities of eMBB and URLLC UEs.
To incorporate UE-AP association as a variable, we define an association matrix $\bm{A}\in\mathbb{R}^{K\times M}$, where $a_{k,m}=1$ indicates association between AP $l$ and UE $k$, and $a_{k,m}=0$ otherwise. Thus, $\mathcal{V}_{k}=\{l:a_{k,m}=1\}$ with cardinality $|\mathcal{V}_{k}|=\sum_{m=1}^{M}a_{k,m}$. Using these definitions, the SINR expression in (5) is reformulated as (6), replacing the summations over subsets $\mathcal{V}_{k}$ with terms involving $\bm{A}$.
@@ IV-A2 Computational Complexity Analysis
$\zeta_{k}\leftarrow w_{k}C_{k}/b_{k}^{\min},\;\forall k$ ;
@@ IV-B UE-AP Association Assignment
Since SINR depends on $\mathbf{A}$ through complex channel combining and interference coupling, P2 is a combinatorial optimization problem that is MINLP problem. Due to the significant computational burden associated with solving MINLP problems optimally (i.e., $\mathcal{O}(2^{KM})$), a heuristic procedure is developed as described below.
@@ V Numerical Analysis
The proposed algorithms are evaluated in a UL CF mMIMO system deployed over a $1\times 1$ km2 area with $M=100$ APs (each having $N=4$ antennas) serving $K$ single-antenna UEs. Both APs and UEs are uniformly distributed, and wrap-around technique [1] is applied to eliminate boundary effects. Following [11], we set $\tau_{p}=10$, $\tau_{c}=200$, $\rho_{p}=\rho_{d}=100~\text{mW/Np}$, and adopt log-normal shadow fading with standard deviation $\sigma_{sh}=8$ dB. To invoke PC more, pilots are randomly assigned to UEs and their power is controlled using [13], while UL data transmission employs open-loop power control. The maximum transmit power for both pilot and data signals is set to 100 mW.
@@ V-1 Comparison of Average Weighted Sum-rate Performance
As expected, the Hybrid scheme achieves the highest weighted sum-rate, outperforming the Proposed scheme by approximately $23\%$, owing to its use of the optimal CVX solver for bandwidth allocation (with greedy fallback in infeasible cases). However, this performance gain comes at the expense of significantly higher computational complexity ($\mathcal{O}(K^{3}+KM)$ compared ot $\mathcal{O}(K\log K+KM\log M)$) and the need for numerical solvers, making the Proposed scheme more suitable for resource-constrained or real-time deployments. However, the Proposed scheme outperforms the Baseline by up to $52\%$. This performance gain is attributed to: (i) the proposed bandwidth allocation (Algorithm 1), which leverages the metric $\zeta_{k}$ capturing UE priority, channel quality, and resource efficiency, and redistributes leftover bandwidth to improve adaptability even when QoS requirements cannot be met, and (ii) the priority-based UE-AP association (Algorithm 2) based on the association potential metric $\Xi_{k,m}$, which complements the per-UE bandwidth allocation. At higher UE densities, the weighted sum-rate performance of all schemes diminish due to increased inter-UE interference and bandwidth constraints.
@@ V-2 Comparison of Average Success Rate Performance
Figs. 2 and 3 illustrate the average success rate, i.e., the fraction of UEs satisfying their respective QoS constraints: minimum-rate requirement (7f) for eMBB and delay constraint (7e) for URLLC UEs. In case of eMBB UEs (Fig. 2), the Proposed scheme achieves $140\%$ and $115\%$ higher success rates than Baseline and Hybrid; for URLLC UEs (Fig. 3), gains are $58\%$ and $43\%$, respectively. Interestingly, while the Hybrid scheme delivers higher weighted sum-rate (Fig. 1), the Proposed scheme excels in QoS satisfaction, demonstrating an effective trade-off. The Hybrid scheme maximizes the aggregate weighted objective by concentrating resources on high-weight UEs, potentially under-serving others below their QoS thresholds. Conversely, the Proposed scheme embeds QoS requirements directly into the $\zeta_{k}$ and $\Xi_{k,m}$ metrics (Algorithms 1 and 2, respectively), factoring in both constraint satisfaction and sum-rate maximization during allocation. This yields a more balanced resource distribution, making Proposed preferable when high success rates are critical despite modest sum-rate sacrifice. These URLLC success rates could be further improved by augmenting the framework with priority-based admission control, granting URLLC UEs preferential admission over eMBB UEs when resources are scarce. Such integration would enable the proposed bandwidth allocation to redistribute resources more effectively, enhancing URLLC reliability while preserving inter-slice fairness. Finally, examining the constraint characteristics, the substantially higher eMBB success rates reflect constraint stringency differences: the heterogeneity in eMBB rate requirements (for premium and standard tiers UEs) allows more flexibility in resource redistribution, whereas URLLC’s uniformly stringent delay constraints offer limited room for adaptation.

### v2

@@ Abstract
This paper addresses the joint optimization of per-user equipment (UE) bandwidth allocation and UE–access point (AP) association to maximize weighted sum-rate while satisfying heterogeneous quality-of-service (QoS) requirements across enhanced mobile broadband (eMBB) and ultra-reliable low-latency communication (URLLC) slices in the uplink of a network slicing-enabled user-centric cell-free (CF) massive multiple-input multiple-output (mMIMO) system. The formulated problem is NP-hard, rendering global optimality computationally intractable. To address this challenge, it is decomposed into two sub-problems, each solved by a computationally efficient heuristic scheme, and jointly optimized through an alternating optimization framework. We then propose (i) a bandwidth allocation scheme that balances UE priority, spectral efficiency, and minimum bandwidth demand under limited resources to ensure fair QoS distribution, and (ii) a priority‑based UE-AP association assignment approach that balances UE service quality with system capacity constraints. Together, these approaches provide a practical and computationally efficient solution for resource‑constrained network slicing scenarios, where QoS feasibility is often violated under dense deployments and limited bandwidth, necessitating graceful degradation and fair QoS preservation rather than solely maximizing the aggregate sum‑rate. Simulation results demonstrate that the proposed scheme achieves up to $52\%$ higher weighted sum-rate, $140\%$ and $58\%$ higher QoS success rates for eMBB and URLLC slices, respectively, while reducing runtime by up to $97\%$ compared to the considered benchmarks.
@@ I Introduction
The sixth-generation (6G) wireless systems must support heterogeneous services with stringent quality-of-service (QoS) demands, primarily enhanced mobile broadband (eMBB) (high-rate, reliable) and ultra-reliable low-latency communication (URLLC) (ultra-low latency, high reliability) [9]. Network slicing (NS) addresses this by partitioning physical infrastructure into logical slices, each optimized for specific QoS objectives [1]. Cell-free (CF) massive multiple-input multiple-output (mMIMO), a promising 6G architecture [11], eliminates cell boundaries by enabling distributed access points (APs) to cooperatively serve user equipments (UEs), achieving superior spectral efficiency (SE), uniform experience, and enhanced reliability through spatial diversity and user-centric connectivity [6]. However, integrating NS into CF mMIMO introduces challenges in resource allocation, UE association, and inter-slice isolation [16, 5].
@@
In the CF mMIMO systems, existing research has predominantly focused on downlink (DL) slicing. For instance, [15] proposes multi-tenant resource allocation maximizing infrastructure revenue while preserving service quality, though without dynamic link selection critical for scalability. The work in [16] introduces a hierarchical slicing framework leveraging multiple time scales to maximize cross-service utility, while [5] exploits spatial diversity to meet heterogeneous URLLC and eMBB requirements under imperfect channel state information. However, they do not address resource allocation in infeasible resource-constrained cases. While uplink (UL) slicing is equally important and poses distinct challenges due to diverse UE demands and limited inter-UE coordination, it has received comparatively less attention in the literature. To date, UL slicing has not yet been explored in the CF mMIMO systems.
@@
Designing UL access is inherently more complex than DL due to a lack of UE coordination and different resource priorities. While DL prioritizes bandwidth and throughput for content delivery, UL emphasizes power management and timely transmission for latency-critical applications. Thus, the stringent QoS requirements imposed by emerging UL-intensive applications (autonomous vehicles, telemedicine, augmented/virtual reality) cannot be met by reusing DL slices, necessitating dedicated UL slicing mechanisms for CF mMIMO systems. However, practical high-density deployments with limited bandwidth frequently violate resource allocation feasibility, where total minimum bandwidth demand exceeds available capacity (further discussed in Section IV-A). In such scenarios, maintaining service continuity for all UEs while ensuring fair QoS distribution becomes critical to prevent complete service outage. This motivates the development of resource allocation strategies that gracefully degrade service quality under scarcity while preserving system-wide fairness and reliability.
@@
In this work, we consider UL slicing in a NS-enabled CF mMIMO framework to support heterogeneous eMBB and URLLC traffic while improving system performance under practical assumptions such as pilot contamination (PC). To the best of our knowledge, the UL slicing problem in the context of the CF mMIMO systems has not yet been addressed in the literature. The main contributions are: (i) we formulate a joint optimization problem for per-UE bandwidth allocation and UE–AP association to maximize weighted sum-rate in NS-enabled UL CF mMIMO systems; due to the NP-hardness of the problem, we propose an alternating-optimization (AO) approach that decomposes it into two tractable sub-problems solved iteratively, (ii) we develop two computationally efficient heuristic algorithms: an efficiency‑oriented bandwidth allocation scheme and a priority‑based UE–AP association scheme, ensuring fair QoS distribution even when feasibility is violated, thereby enabling graceful degradation under resource scarcity, and (iii) extensive numerical results demonstrate that the proposed solution achieves a favorable trade-off between weighted sum-rate and QoS satisfaction while substantially lowering computational cost compared to an interior-point solver-based solution.
@@ II-C Channel Estimation Model
The UL transmission consists of two main stages: training and data transmission. In each coherence interval of length $\tau_{c}$ symbols, $\tau_{p}$ symbols are allocated for channel estimation and the rest is for data transmission. During UL training, all UEs transmit their assigned pilot sequences, which are received by the APs for channel estimation. As in practical systems, we assume that $K\gg\tau_{p}$, implying that some UEs reuse the same pilot sequences, leading to PC. The channel is estimated using the minimum mean square error (MMSE) approach and the quality of the estimated channel $\hat{\mathbf{g}}_{k,m}$ is quantified by the mean-square of the estimate: $\gamma_{k,m}\triangleq\mathbb{E}\left[\|\hat{\mathbf{g}}_{k,m}\|^{2}\right]=\sqrt{\tau_{p}\rho_{p}\,\eta_{k}^{p}\,}\beta_{k,m}\,c_{k,m}$. The MMSE scaling coefficient $c_{k,m}$ is given by
@@ II-D Uplink Data Transmission Model
During UL transmission, each UE $k\in\mathcal{K}$ sends data symbols to its cooperating APs, [^1] Each AP receives all UEs’ signals but decodes only those from its associated UEs according to the UE-AP association scheme. which forward them to the CPU for joint decoding using channel estimates. URLLC UEs, having short packets, are modeled using finite block length theory rather than the Shannon capacity [16]. The achievable rate expression for both eMBB and URLLC UEs is given at the top of the next page, where $V_{k}=1-(1+\text{SINR}_{k})^{-2}$ is the channel dispersion, ${Q}^{-1}(\cdot)$ is the inverse Gaussian Q-function, $\theta$ is the decoding error probability, and $L_{k}$ is the packet size of UE $k$. The closed-form signal-to-interference-plus-noise ratio (SINR) expression for UE $k$ ($\text{SINR}_{k}$) is written as (5) using similar steps described in [7], where $\eta_{k}^{d}\in(0,1]$ denotes the data power control coefficient of UE $k$ and $\rho_{d}$ is the normalized UL data SNR.
@@ II-E Delay Model
To capture the latency characteristics of delay-sensitive services such as URLLC, we adopt an $M/M/1$ queuing model as used in [14] to characterize the UL queuing and transmission delay experienced by each UE. For UE $k\in\mathcal{K}_{\mathrm{URLLC}}$, let $\lambda_{k}$ denote the average packet arrival rate (in packets per second). The service rate of UE $k$ is $\mu_{k}=R_{k}/L_{k}$ (packet size in bytes). So, under the stability condition $\mu_{k}>\lambda_{k}$ [8], the total delay experienced by UE $k$, denoted by $D_{k}$, comprises both queuing and transmission delays and is expressed as: $D_{k}=\frac{1}{\mu_{k}-\lambda_{k}}=\frac{1}{\frac{R_{k}}{L_{k}}-\lambda_{k}},\quad\forall k\in\mathcal{K}_{\mathrm{URLLC}}$.
To satisfy the QoS requirements of latency-critical applications, the total delay must not exceed a predefined maximum tolerable latency threshold $D_{k}^{\mathrm{max}}$. This requirement imposes the following constraint on the achievable UL rate: $\frac{R_{k}}{L_{k}}\geq\lambda_{k}+\frac{1}{D_{k}^{\mathrm{max}}},\quad\forall k\in\mathcal{K}_{\mathrm{URLLC}}$. This constraint ensures that the combined queuing and transmission delays remain within acceptable bounds for each UE of $\mathcal{K}_{\mathrm{URLLC}}$, thereby preserving the reliability and responsiveness required by mission-critical services.
@@ III Problem Formulation
Our objective is to jointly optimize the per-UE bandwidth allocation and UE-AP association assignment for maximizing the weighted sum-rate performance under fixed slice-wise bandwidth $B_{s},\ s\in\mathcal{S}$, while satisfying service-specific QoS requirements. The weights $w_{k}$ reflect heterogeneous service priorities of eMBB and URLLC UEs.
To incorporate UE-AP association as a variable, we define an association matrix $\mathbf{A}\in\{0,1\}^{K\times M}$, where $a_{k,m}=1$ indicates association between AP $m$ and UE $k$, and $a_{k,m}=0$ otherwise. Thus, $\mathcal{V}_{k}=\{m:a_{k,m}=1\}$ with cardinality $|\mathcal{V}_{k}|=\sum_{m=1}^{M}a_{k,m}$. Using these definitions, the SINR expression in (5) is reformulated as (6), replacing the summations over subsets $\mathcal{V}_{k}$ with terms involving $\mathbf{A}$.
@@ IV-A2 Computational Complexity Analysis
$\zeta_{k}\leftarrow w_{k}\mathrm{SE}_{k}/b_{k}^{\min},\;\forall k$ ;
@@ IV-B UE-AP Association Assignment
Since SINR depends on $\mathbf{A}$ through complex channel combining and interference coupling, P2 is a combinatorial optimization problem that is an MINLP problem. Due to the significant computational burden associated with solving MINLP problems optimally (i.e., $\mathcal{O}(2^{KM})$), a heuristic procedure is developed as described below.
@@ V Numerical Analysis
The proposed algorithms are evaluated in a UL CF mMIMO system deployed over a $1\times 1$ km2 area with $M=100$ APs (each having $N=4$ antennas) serving $K$ single-antenna UEs. Both APs and UEs are uniformly distributed, and the wrap-around technique [2] is applied to eliminate boundary effects. Following [11], we set $\tau_{p}=10$, $\tau_{c}=200$, $\rho_{p}=\rho_{d}=100~\text{mW/Np}$, and adopt log-normal shadow fading with standard deviation $\sigma_{sh}=8$ dB. To induce PC effects more, pilots are randomly assigned to UEs and their power is controlled using [13], while UL data transmission employs open-loop power control. The maximum transmit power for both pilot and data signals is set to 100 mW.
@@ V-1 Comparison of Average Weighted Sum-rate Performance
As expected, the Hybrid scheme achieves the highest weighted sum-rate, outperforming the Proposed scheme by approximately $23\%$, owing to its use of the optimal CVX solver for bandwidth allocation (with greedy fallback in infeasible cases). However, this performance gain comes at the expense of significantly higher computational complexity ($\mathcal{O}(K^{3}+KM)$ compared to $\mathcal{O}(K\log K+KM\log M)$) and the need for numerical solvers, making the Proposed scheme more suitable for resource-constrained or real-time deployments. However, the Proposed scheme outperforms the Baseline by up to $52\%$. This performance gain is attributed to: (i) the proposed bandwidth allocation (Algorithm 1), which leverages the metric $\zeta_{k}$ capturing UE priority, channel quality, and resource efficiency, and redistributes leftover bandwidth to improve adaptability even when QoS requirements cannot be met, and (ii) the priority-based UE-AP association (Algorithm 2) based on the association potential metric $\Xi_{k,m}$, which complements the per-UE bandwidth allocation. At higher UE densities, the weighted sum-rate performance of all schemes diminishes due to increased inter-UE interference and bandwidth constraints.
@@ V-2 Comparison of Average Success Rate Performance
Figs. 2 and 3 illustrate the average success rate, i.e., the fraction of UEs satisfying their respective QoS constraints: minimum-rate requirement (7f) for eMBB and delay constraint (7e) for URLLC UEs. In case of eMBB UEs (Fig. 2), the Proposed scheme achieves $140\%$ and $115\%$ higher success rates than Baseline and Hybrid; for URLLC UEs (Fig. 3), gains are $58\%$ and $43\%$, respectively. Interestingly, while the Hybrid scheme delivers higher weighted sum-rate (Fig. 1), the Proposed scheme excels in QoS satisfaction, demonstrating an effective trade-off. The Hybrid scheme maximizes the aggregate weighted objective by concentrating resources on high-weight UEs, potentially under-serving others below their QoS thresholds. Conversely, the Proposed scheme embeds QoS requirements directly into the $\zeta_{k}$ and $\Xi_{k,m}$ metrics (Algorithms 1 and 2, respectively), factoring in both constraint satisfaction and sum-rate maximization during allocation. This yields a more balanced resource distribution, making the Proposed scheme preferable when high success rates are critical despite modest sum-rate sacrifice. These URLLC success rates could be further improved by augmenting the framework with priority-based admission control, granting URLLC UEs preferential admission over eMBB UEs when resources are scarce. Such integration would enable the proposed bandwidth allocation to redistribute resources more effectively, enhancing URLLC reliability while preserving inter-slice fairness. Finally, examining the constraint characteristics, the substantially higher eMBB success rates reflect constraint stringency differences: the heterogeneity in eMBB rate requirements (for premium- and standard-tiers UEs) allows more flexibility in resource redistribution, whereas URLLC’s uniformly stringent delay constraints offer limited room for adaptation.